Systems Staff Meeting Notes
July 10, 2006
July 10, 2006
Present: Cynthia Bowker, Will Chaney, Suzy Covey, Debra Fetzer, Winston Harris, Bobbie Meng, Tom Minton, Chris Nicolich, Geoffrey Stewart, Ying Tang, Stephen WilliamsAnnouncements
Ongoing Systems Updates
- Library Council meeting last week - new Library-wide policy on cooking. There are only three locations where food can be cooked: 5th floor West staff lounge, 1st floor MSL staff lounge and 3rd floor East staff lounge. No microwaves or toaster allowed within departments. Eating at your desk is still allowed, however. Coffee pots are okay. Formal announcement should go out in a few days.
- Budget is flat. No more money than last year.
- Director of Library search - starting to get underway.
- Customer Service workshop upcoming - will be mandatory.
- Staff Structure Reorg Report - available on the HR Web site. Will is trying to make sure we are eventually included in the study. We'd like to be making comparable salaries in comparison to the rest of campus IT. Need to look at both sides of the coin.
- OPS budget - need to advise Business Services if we think it is going to change for next year.
- Systems temporary move is getting close. Rob is going to bring some boxes by for packing. Carpeting issue: Bobbie is sensitive to something in the carpeting in West.
- Judaica is moving into both West and East. The collection and one person will be housed in West while one person will be in East.
- Amy Polk should be starting this Friday, July 14.
- Ordering restarted.
- There is a lot for West that needs to be ordered. Currently waiting for a Dell quote for classroom and media room installation.
- The old Sun server in Chris' office can be surveyed out. Need to format the HD.
- We will employ liaisons within a week or so to help set up the high-end media area. There is a long list of software to be installed on these machines. Will need to do it one machine at a time.
Software Projects Status
- Installation of keyboard, mouse, monitor, UPS in server room - still waiting for Facilities to get the power set up.
- Backup for Greenstone Development server - Geoffrey researched - the latest extensions with the right drives & drives should be on the CD. We need to upgrade the version that is on Greenstone - make sure it includes extensions.
- Training database - will record who has conducted and attended trainings. The software folks are working on this.
- Quality Control - standard for testing before programs are released. Will write something up so that the party has to sign-off on acceptance.
- Grant Tracking database - still in review.
- ILLIAD - went to OCLC control about a week and a half ago. There are still some issues with our software. We did a pretty good job - just ironing out the last few items. They wanted our email server to route their email - just relay them. It did not work. Turns out, the OCLC virus checker was blocking port 45. Once everything is worked out, the ILLIAD server will be taken down.
- ERES interface - out for approval. Reviews positive. Issue of authentication - working on it.
- Ying - finishing up two things for Greenstone/UFDC, then she will be doing documentation.
- Circulation manual - up and available online.
- MetaLib is up and running for the rest of the summer. It will be used and we hope to get feedback about it.
- Blog server issue - we have a test version of the software on Tom and Jason's computers.
- Meeting held.